OpenStudy (anonymous):

Solve for x: 3(x + 1) = -2(x - 1) + 6. 1 4 5 25

OpenStudy (shamil98):

distributive property. 3x + 3 = -2x + 2 + 6 3x + 3 = -2x + 8 5x = 5 x = 1